A harmonically moving transverse wave on a string has a maximum particle velocity and acceleration of 3 m/s and 9 m/s2 respectively. Velocity of wave is 20 m/s , then the wave form is

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
maximum particle velocity =wA=3
acceleration max=w2A=9
w=3
A=1

Vwave=20

f=w/2π=3/2π
v=lf
=>l=v/f
=>l=20x2π/3
=>l=40π/3

k=2π/l
=>k=6/40

The equation of the wave form is= Asin(wt-kx)
=sin(3t-6x/40).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)